How to make this factorized quadratic equation into an standard form of an quadratic equation?
(x-1)(x-2) = 0
Thank you!:)

Answer by mananth(16946)

(x-1)(x-2) = 0
x(x-2)-1(x-2)=0
x^2-2x-x+2=0
add like terms
x^2-3x+2=0